OpenStudy (anonymous):

I'm having trouble with this definite integral. a=1/48 b=1/16 csc(8πt) cot(8πt) dt

OpenStudy (anonymous):

\[\int\limits_{1/48}^{1/16}\csc(8πt)*\cot(8πt)*dt\]

OpenStudy (anonymous):

anybody......

OpenStudy (anonymous):

I did every one of them from my homework without any trouble, but can't solve this one... So if I make u=8*pi*t du=8pi*dt (du/8pi)=dt so: \[\int\limits_{1/48}^{1/16} \csc(u)*\cot(u)*(du/8π)\]

OpenStudy (mathteacher1729):

Hint, the derivative of csc(u) = -cot(u)csc(u)

OpenStudy (anonymous):

Aren't I supposed to take the antiderivative?

OpenStudy (anonymous):

Oh wait, I think i see.... Gimme a sec, don't go anywhere please!

OpenStudy (mathteacher1729):

You're on the right path, but remember to change your limits! u = 8 pi t so u(1/16) = 8 pi (1/16) = pi/2 and u(1/48) = 8 pi (1/48) = ... etc. THEN you can go and do the rest of the integral, which should be easy from here on in. :)

OpenStudy (anonymous):

Yeah I see that^^, Thanks, but I can't just plug those numbers into csc(u)*cot(u)*(du/8pi), right? I have to plug those upper and lower limits into the antiderivative of csc(u)*cot(u)*(du/8pi), which is what i am struggling with.

OpenStudy (anonymous):

Oh my gosh, I'm such an idiot. i see. Now your hint makes sense to me!

OpenStudy (anonymous):

f(x)= -csc(x)+(x/8π) + C f'(x)= csc(x) * cot(x) * (dx/8pi) If my calculations are correct.

OpenStudy (mathteacher1729):

Not quite... if f(x) = csc(u) then f'(x) = -cot(u)*csc(u)

OpenStudy (mathteacher1729):

And you already reduced your integral to integral of 1/(8pi) cot(u)*csc(u) du

OpenStudy (mathteacher1729):

So you're like... right on the doorstep of this.

OpenStudy (anonymous):

Now i'm just going to plug in u(1/46) and u(1/16) into -csc(u)+((u/8pi)???

OpenStudy (anonymous):

So now do I plug in the upper and lower limits (u(1/16) and u(1/46)) into -csc(u)+((u/8pi) and find the difference????

OpenStudy (anonymous):

I can't figure it out! damnit

OpenStudy (anonymous):

[-csc(pi/6)+((pi/6)/(8pi)) - -csc(pi/2)+((pi/2)/(8pi))] = [-15/16 - -95/48] = 25/24 Which IS NOT RIGHT......... Okay, F*ck this, I've spent two hours on this and i have other sh*t to do...